OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − x2 = −11x + 24? (4 points) (−3, 0) and (8, 0) (3, 0) and (−8, 0) (−3, 0) and (−8, 0) (3, 0) and (8, 0)

OpenStudy (anonymous):

please!!!!!

OpenStudy (anonymous):

ASAP

OpenStudy (sburchette):

adding x^2 to both sides we get\[y=x^2-11x+24\]To find the x-intercepts of a function, substitute 0 for y and then solve for x. Because it is a quadratic function, you can expect 2 intercepts unless they are complex roots. Substituting 0 for y we have\[0=x^2-11x+24\] Now we need to factor the right side. We need to find two values such that when they are multiplied they equal 24 and when they are added they equal -11. -8*-3=24 and -8-3=-11 so we know that it will factor into \[0=(x-3)(x-8)\]Now we can set each factor equal to zero to obtain the solutions which are\[x=3,x=8 \]

OpenStudy (anonymous):

also I need help with this one What are the solutions of 6x2 − x − 15 = 0? (4 points) x = , x = −3 x = −, x = 3 x = , x = − x = −, x =